OpenStudy (briana.img):

The perimeter of a triangular playground is 270 feet. If the ratios of the measures of the sides of the triangle are 5:6:7, find the measure of the sides of the playground.

OpenStudy (briana.img):

A.60 feet, 80 feet, 100 feet B .75 feet, 90 feet, 105 feet C. 50 feet, 60 feet, 70 feet D. 48 feet, 72 feet, 144 feet

OpenStudy (anonymous):

tricky

OpenStudy (briana.img):

yeah i thought you just had to divide but that doesn't seem to work..

OpenStudy (anonymous):

i hate math lol

OpenStudy (briana.img):

me too LOL retaking two classes

OpenStudy (anonymous):

ouch </3

OpenStudy (anonymous):

please don't kill me if I'm wrong

OpenStudy (briana.img):

nah it's okay

OpenStudy (anonymous):

lets go with B

OpenStudy (briana.img):

okay well i'm trying to look for an example in my lesson to see if i could figure it out that way, but if i can't i'll just go with b

OpenStudy (anonymous):

first fan1 oh yeah you're welcome

OpenStudy (briana.img):

found an example :p

OpenStudy (anonymous):

yay!

OpenStudy (briana.img):

why are you going with b? are you just guessing? lol

OpenStudy (anonymous):

i tried

OpenStudy (briana.img):

ooo ok

OpenStudy (briana.img):

i put that as the answer, i have a few more questions to get through before the test is done so i'll let you know if it's right

OpenStudy (cwrw238):

change the ratios to fractions 5:6:7 5+6+7 = 18 so we we have 5/18 , 6/18 and 7/18 or 5/15 /1/3 and 7/18 so multiply 270 by each of these fractions to get your lengths

OpenStudy (johnweldon1993):

Hmm well lets see..the perimeter is 270 right? Assume we just use the 3 4 and 5 provided \[\large 5^2 + 6^2 = 7^2\] It works out so lets add them \[\large 5 + 6 + 7 = 18\] Now the perimeter we need is \(\large \frac{270}{18} = 15\) times bigger than that perimeter, so lets multiply each side we have (5 6 and the 7) by that 15 \[\large 5 \times 15 = 75\] \[\large 6 \times 15 = 90\] \[\large 7 \times 15 = 105\]

OpenStudy (cwrw238):

note the middle value is 270 * 1/3 = 90

OpenStudy (briana.img):

@johnweldon1993 Thank you for replying to my question! I clearly understood how you solved it and finally understand how to do it!

OpenStudy (cwrw238):

* sorry it should read 5/18, 1/3 and 7/18

OpenStudy (johnweldon1993):

You are very welcome! :)